OpenStudy (anonymous):

Naomi bikes the 40 miles to Hillsboro averaging a certain speed. The return trip is made at a speed 6 miles per hour slower. The total time for the round trip is 14 hours. Find Naomi's average speed on each part of the trip.

OpenStudy (yttrium):

Let x be the initial speed. x-6 be the return trip speed y be the amount of initial time 14 - y be the the return trip time Since Naomi traveled the same distance because she only had a return trip you can formulate d=d. Such that d = St. Where d is the distance, s is the seed and t is the time. Can you do the next steps now? Or still needs help?

OpenStudy (anonymous):

Could you do the next step please?

OpenStudy (yttrium):

*i mean speed

OpenStudy (yttrium):

Wait I think something is wrong here! hahahah!

OpenStudy (jack1):



OpenStudy (jack1):

speed there = x miles/hr speed back = y miles per hour note: the return speed is 6 miles per hour slower than the speed there so y = x - 6

OpenStudy (jack1):

total time = 14hrs total distance = 80 miles average speed would be 80 miles / 14 hrs = about 5.7 miles per hours but actual speed (x and y) ...

OpenStudy (jack1):

are 80 miles in 14 hrs 40 miles / x (miles per hour) there = time to get there 40 miles / y (miles per hour) back = time to get home as y = x - 6 40 miles / (x - 6) = time to get home time to get there + time to get home = 14 hrs solve for x

OpenStudy (jack1):

so what are your 2 possible answers for x? @justine97 ?

OpenStudy (anonymous):

I'm writing the problem right now, just give me a sec

OpenStudy (jack1):

no worries, sorry

OpenStudy (anonymous):

So I got 40/x + 40/x-6 = 14

OpenStudy (anonymous):

But I'm having trouble solving for x

OpenStudy (jack1):

\[\frac{ 40 }{ x } + \frac{ 40 }{x - 6 } = 14\] cross multiply to get common denominator \[\frac{ 40(x-6) }{ x(x - 6) } + \frac{ 40x }{ x(x-6) } = 14\] common base, so simply add the numerators together (top line) \[\frac{ 40(x-6) + 40x }{ x(x-6) } = 14\] add like terms, expand brackets and simplify \[\frac{ 40x - 240 + 40x }{ x^2 -6x } = 14\] \[\frac{ 80x - 240 }{ x^2 - 6x } = 14\] times both sides by (x^2 - 6x) to remove denominator

OpenStudy (jack1):

\[80x - 240 = 14(x^2 -6x)\] \[80x - 240 = 14x^2 - 84x\] \[0 = 14x^2 - 164x + 240\] solve using quadratic formula

OpenStudy (jack1):

ur familiar with this one... yeah?

OpenStudy (anonymous):

Okay, got it! :)

OpenStudy (anonymous):

2(7x-12)(x-10) = 0

OpenStudy (jack1):

in this case: a = 14 b = -164 c = 240 so solve using the formula and u should get something close to: x = 10 or x = 1.7 now 1.7 is a crazy slow speed to be travelling as to get 40 miles at 1.7 miles per hour would take 23 ish hours, which is heaps more that the 14 hours total to travel 80 miles ... but at 10 miles per hour, it would take 4 hours to do the first leg... which is sensible

OpenStudy (anonymous):

So we throw out 1.7 because the number doesn't make sense?

OpenStudy (jack1):

and given the return trip is 6 miles slower 10 - 6 = 4 so y = 4 miles per hour to travel 40 miles would take... 10 hours so 4 hrs there, 10 hours return: 14 hours total... sweet

OpenStudy (jack1):

we throw out the 1.7 as it dosent fit our equation, we always do a logic test on the answers as there's 2 possible answers to every sqr root (+ve and -ve)

OpenStudy (jack1):

ie 4^2 = 16 but (-4)^2 also = 16

OpenStudy (anonymous):

WOW, this is amazing! Thank you so so so much! :)

OpenStudy (jack1):

nah all good dude, happy to help hey it's all makin sense now tho?

OpenStudy (anonymous):

Yes! This made a lot of sense. Again, you were super helpful and patient. I appreciate the help!

OpenStudy (jack1):

anytime hey, slaters ;D